# Governments Should Mandate Tiered Anonymity on Social-Media Platforms to Counter Deepfakes and LLM-Driven Mass Misinformation

### v1

@@ Abstract
This position paper argues that governments should mandate a three-tier anonymity framework on social-media platforms as a reactionary measure prompted by the ease-of-production of deepfakes and large-language-model-driven misinformation. The tiers are determined by a given user’s reach score: Tier 1 permits full pseudonymity for smaller accounts, preserving everyday privacy; Tier 2 requires private legal-identity linkage for accounts with some influence, reinstating real-world accountability at moderate reach; Tier 3 would require per-post, independent, ML-assisted fact-checking, review for accounts that would traditionally be classed as sources-of-mass-information.
@@ 1 Introduction
Online anonymity was originally a shield for ordinary speakers, political dissidents, and vulnerable groups. However, when algorithmic amplification gives a single post the reach of a broadcaster, blanket anonymity becomes a public-safety liability. We therefore argue that identity obligations should scale with influence.
Our proposed three-tier model, summarized in Table 1, assigns obligations by some notion of reach (e.g. a weighted sum of followers, shares, views, etc.). This is further explored in Section 5. Tier 1 preserves full pseudonymity for low-reach accounts; Tier 2 requires a platform-held legal-identity link once a predefined reach threshold is crossed; Tier 3 adds independent, ML-assisted fact-checking for mass-reach content.
Table 1: Proposed tiered anonymity framework. Tier thresholds are discussed in Section 5, are generally illustrative, and should be calibrated per-platform. A single post that crosses a threshold retroactively elevates the account to the corresponding tier.
@@
We ground our proposal in using empirical evidence from Reddit’s community moderation approach. Volunteer moderators already converge on proportional governance: as subreddit traffic grows, moderators introduce karma [^1] In Reddit’s context, this denotes an aggregate reputation metric equal to the net difference between positive and negative votes that a user’s submissions and comments receive. It thus functions as a quantifiable proxy for community trust and is frequently employed as an eligibility threshold for posting or moderation privileges. minimums, manual reviews, or identity checks before posts appear [6, 7, 8]. These organic practices demonstrate operational feasibility and indicate latent demand for tiered accountability that transcends any single platform architecture.
We further show that there is a viable regulatory pathway to achieving our proposal. The European Union’s Digital Services Act already requires marketplaces to verify business users and offers a blueprint for identity-linked content duties [9]. The UK Online Safety Act obliges “Category 1” [^2] Typically, major social media platforms services to give adults tools for filtering anonymous accounts and mandates that platforms offer identity verification [10]. By drawing on these precedents, legislators can embed tiered anonymity into safety models and ranking systems without prohibiting pseudonymity outright.
@@
    We chart a concrete, jurisdiction-spanning regulatory pathway that leverages existing DSA and Online Safety Act provisions to operationalize the model.
@@ 3 Friction, Identity, and Accountability
The starting point for any meaningful reform of online anonymity must confront a central tension in liberal democracies: the commitment to free expression versus the need to mitigate its weaponization. Classic accounts of speech rights, from John Stuart Mill to modern First Amendment jurisprudence, treat expression as a public good – presumptively beneficial and self-regulating [21, 22, 23]. Yet in algorithmically mediated environments, where LLMs can generate plausible falsehoods at scale and deepfakes blur the boundary between reality and fabrication, virality becomes decoupled from both truth and reputation, and the foundational assumptions underpinning these traditions begin to unravel.
Friction – in the form of verification, moderation, or traceability – is often framed as a threat to openness [5]. That said, friction is a democratic design feature [24]. In physical communities, social friction arises from reputational consequences, geographic co-presence, and mutual visibility. One is less likely to spread inflammatory falsehoods in a town hall than online, not because one is more moral, but because the social costs are real and immediate. Digital platforms, in contrast, systematically dissolve these frictions. Recommender systems prioritize engagement, not deliberation; speed trumps reflection; and pseudonymity attenuates accountability [25].
@@
We argue that identity obligations must scale with content reach , particularly to mitigate the systemic risks posed by synthetic media as described in Section 2. This is not a blanket call for real-name policies, which have been rightly criticized for silencing vulnerable speakers [30, 31]. Instead, it is a call for proportional identity calibration [32], wherein pseudonymity is preserved for low-reach users, while higher-tier actors must submit to private identity verification and, ultimately, to structured content review [33]. This approach mirrors how democratic institutions already manage power: with increasing transparency and accountability as influence grows [34].
Our Reddit case study, described in Section 4, illustrates this principle in practice. As communities expand in size and influence, moderation architectures evolve from permissive to hierarchical: identity checks, posting restrictions, and content approvals become the norm. These organically emergent structures reflect a collective intuition: that scale demands scrutiny and visibility must be earned.
@@ 4 Reddit Case Study in Community Moderation at Scale
Quantitative work finds a positive correlation between subreddit size and the likelihood of (i) entry gates (minimum account age/karma) [8], (ii) pre-publication queues [8, 41], and (iii) ex post identity checks [7]. In other words, moderators intuitively impose proportional friction: low-reach users post freely; higher-reach content encounters verification or review. Reddit thus supplies real-world evidence that tiered anonymity is operationally feasible and socially accepted when the costs of influence are borne chiefly by those who wield it.
Table 2: Identity and moderation norms on major social-media platforms as of May 2025. “Tiered” denotes any mechanism in which obligations or scrutiny escalate with audience size or monetization status.
@@
Additional scrutiny for $100$k+ channels
@@
Competing platforms provide no comparable venue for community-level rule-making. Facebook real-name enforcement, X’s paid “blue check”, and YouTube’s purely algorithmic filters all exemplify top-down moderation with minimal local discretion or tiering. Comparative studies confirm that Reddit alone relies “more or less on self-moderation by volunteers”, producing a distinctive, multi-layer oversight regime [42]. We summarize our findings regarding identity and moderation norms in all current major social media platforms in Table 2. YouTube provides the closest analogue to our tiered system: “new” channels face automated checks, Partner-Program creators add identity and monetization audits, and six-figure-subscriber channels receive further manual review and provenance badges.
@@ 5 Proposed High-level Technical Implementation
Most large platforms already store granular engagement telemetry (followers, impressions, reshares, watch-time). A platform-side service can aggregate these signals into a rolling reach score and map it to the tier thresholds proposed in Table 1. To avoid the effect of one-off viral spikes, thresholds should be evaluated over some time period, say a three-month moving window, and updated nightly. When a score first crosses a threshold, a workflow flags the account for tier elevation and temporarily rate-limits outbound posts until the verification step – ID upload for Tier 2; fact-checking for Tier 3 – is completed.
[…]
@@ 6 Current Legal Precedents and Regulatory Infrastructure
6.1 European Union: From KYBC to Reach-Based Identity Accountability
@@
More significantly, Articles 34 and 35 impose systemic risk obligations on Very Large Online Platforms (VLOPs) – defined by a monthly audience threshold – explicitly linking reach to responsibility [48]. These risk obligations are particularly salient in the age of generative AI, again linking back to the discussion in Section 2. This capacity decouples virality from veracity, amplifying the urgency of proportional identity and transparency duties for high-reach users. This sets a critical precedent: the broader a user or platform’s influence, the greater the required diligence. Article 9 further enables identity disclosure in response to illegality, reinforcing a principle of proportionality that mirrors the core logic of tiered anonymity [49].
Moreover, complementary frameworks like the AI Act and proposed AI Liability Directive further strengthen this trajectory. By requiring labeling of synthetic media and audit trails for AI systems, the EU is already enforcing traceability in high-risk communicative environments [50, 51]. Our proposed tiered framework fits squarely within this expanding digital acquis. These instruments, taken together, suggest that scalable identity obligations based on content reach are not only compatible with EU law – they are its logical extension.
@@ 6.2 United Kingdom: The Online Safety Act and Voluntary Verification
In contrast, the UK’s Online Safety Act 2023 [10] establishes a statutory duty of care on digital platforms, particularly those classified as Category 1 services – platforms with significant reach and functionality. Under the accompanying Categorization of Regulated Services Threshold Conditions Regulations 2024 [52], these platforms are required to offer adult users the option to verify their identity and to provide tools enabling content filtering based on verification status. This framework introduces a layered reputational infrastructure while preserving the right to anonymity, laying the conceptual groundwork for our proposed tiered identity regime.
However, this identity framework remains voluntary and reputational rather than mandatory and enforceable. Users may choose to verify themselves, and others may opt to filter content accordingly; but no binding obligations are imposed on high-reach users who remain anonymous. Legal mechanisms such as the Norwich Pharmacal Orders and the UK–US CLOUD Act [53, 54] already allow for identity disclosure under judicial or governmental request, affirming that anonymity online is not absolute but subject to contextual limits.
Nevertheless, the current UK regulatory landscape lacks a proactive mechanism linking user influence – measured by visibility, engagement, or monetization – to identity obligations. We argue that this omission is increasingly untenable in an era of algorithmic virality where individuals can rapidly attain significant reach with little to no accountability.
A logical evolution of the Online Safety Act would be to mandate identity verification for users who exceed a defined influence threshold, as discussed in Section 5. Such a reform would convert identity verification from a reputational indicator into a mechanism of enforceable accountability. By embedding this obligation within the existing statutory framework, the UK could pioneer a rights-preserving yet responsibility-tiered model of online governance – one that maintains anonymity for everyday users while ensuring that high-reach actors meet proportionate standards of transparency and legal traceability.
@@ 6.3 United States: First-Amendment Boundaries and Conditional Immunity
The United States presents the most challenging jurisdiction for any form of compelled identity regulation due to robust First Amendment protections and the shield of Section 230 of the Communications Decency Act [55]. American courts have repeatedly upheld the right to anonymous speech, particularly in digital spaces. Landmark cases such as Doe v. Cahill and Dendrite Int’l, Inc. v. Doe No. 3 [56] require plaintiffs seeking to unmask anonymous users to meet stringent standards, such as presenting a prima facie case of harm and passing a balancing test that weighs the speaker’s right to anonymity.
Despite this, momentum is growing at the federal level toward rethinking the blanket nature of Section 230 immunity. Legislative proposals, including bipartisan efforts, have increasingly considered conditioning immunity on a platform’s compliance with transparency and good-faith content moderation practices [57, 58]. Rather than mandating identity disclosure, these proposals suggest a path for indirect, incentive-based regulation that respects constitutional limits while introducing mechanisms of accountability.
In this context, our proposed tiered identity approach offers a legally viable and technically feasible approach. Platforms could retain full Section 230 protections only if they adopt a structured system of user obligations based on influence. Such a framework would allow users to remain anonymous at lower tiers but require incremental disclosures or review processes as their reach (and thus potential for public impact) increases. Imposing obligations via a user’s reach or other platform-side metrics such as engagement telemetry and monetization enrollment allows the system to remain content-neutral, which is crucial for surviving constitutional scrutiny [59, 60].
@@ 8 The Global Momentum for Conditional Pseudonymity
The international policy environment is already increasingly converging around the idea that identity obligations should scale with user influence. Early efforts to regulate anonymity include South Korea’s real-name verification law, in effect from 2009 to 2012, that sought to impose identity disclosure universally. That approach proved both legally unsustainable and practically ineffective. The Korean Constitutional Court invalidated the policy for violating freedom of expression, and subsequent research showed it failed to reduce online harms in any measurable way [66, 67]. The lesson was clear: blanket identity mandates are blunt instruments that overreach without precision.
Since then, regulatory energy has shifted toward more granular, influence-sensitive models. In India, the 2023 Draft Digital India Bill introduces a risk-based classification framework for digital intermediaries, suggesting a shift toward more nuanced regulatory obligations based on the type and scale of service; without explicitly extending these obligations to individual users or calibrating them to user influence [68]. Australia’s eSafety Commissioner has advanced similar proposals, calling for the traceability of high impact accounts, particularly those linked to harmful or AI-generated content [69]. Meanwhile, the European Commission has initiated consultations on “influence transparency”, exploring how verification requirements might apply to accounts disseminating politically sensitive or synthetic media [9, 70]
@@
9 Cross-Jurisdictional Implementation and Extraterritorial Reach
Implementing a tiered anonymity framework in a globally interconnected internet ecosystem presents significant enforcement challenges. While Reddit shows that moderation hierarchies can emerge organically, its reliance on volunteer governance is difficult to replicate on more commercial platforms like Meta, YouTube, or X. These platforms operate across multiple jurisdictions but often default to the legal norms of their home country – typically the United States – resulting in fragmented regulatory oversight [75].
To scale tiered anonymity, enforcement must be institutional, driven by governments and platforms rather than individual users. Governments in regions such as the EU, UK, and US already exercise regulatory authority over platforms operating within their borders. This authority can be extended extraterritorially, as seen with the EU’s General Data Protection Regulation (GDPR), which extends obligations beyond EU borders through data adequacy requirements and reputational enforcement mechanisms [76, 77].

### v2

# Tiered Anonymity on Social-Media Platforms as a Countermeasure against Deepfakes and LLM-Driven Mass Misinformation

@@ Abstract
We argue that governments should mandate a three-tier anonymity framework on social-media platforms as a reactionary measure prompted by the ease-of-production of deepfakes and large-language-model-driven misinformation. The tiers are determined by a given user’s reach score: Tier 1 permits full pseudonymity for smaller accounts, preserving everyday privacy; Tier 2 requires private legal-identity linkage for accounts with some influence, reinstating real-world accountability at moderate reach; Tier 3 would require per-post, independent, ML-assisted fact-checking, review for accounts that would traditionally be classed as sources-of-mass-information.
@@ 1 Introduction
Online anonymity began as one of the internet’s core protective mechanisms, allowing ordinary people, dissidents, and vulnerable groups to speak freely without fear of retaliation. However, when algorithmic amplification gives a single post the reach of a broadcaster, blanket anonymity becomes a public-safety liability. We therefore argue that identity obligations should scale with influence. Crucially, the aim is not to censor speech but to regulate its amplification and reach, preserving free expression while curbing the outsized impact of unverified high-reach accounts.
Our proposed three-tier model, summarized in Table˜1, assigns obligations by some notion of reach (e.g. a weighted sum of followers, shares, views, etc.). Tier 1 preserves full pseudonymity [^1] We use “pseudonymity” to mean the use of persistent screen names not linked to offline identity for low-reach accounts; Tier 2 requires a platform-held legal-identity link once a predefined reach threshold is crossed; Tier 3 adds independent, ML-assisted fact-checking for mass-reach content.
Table 1: Proposed tiered anonymity framework. Tier thresholds are discussed in Section˜5, are generally illustrative, and should be calibrated per-platform. A single post that crosses a threshold retroactively elevates the account to the corresponding tier.
@@
We ground our proposal in using empirical evidence from Reddit’s community moderation approach. Volunteer moderators already converge on proportional governance: as subreddit traffic grows, moderators introduce karma [^2] In Reddit’s context, this denotes an aggregate reputation metric equal to the net difference between positive and negative votes that a user’s submissions and comments receive. It thus functions as a quantifiable proxy for community trust and is frequently employed as an eligibility threshold for posting or moderation privileges minimums, manual reviews, or identity checks before posts appear [50, 72, 31]. This means baseline pseudonymity for ordinary users is largely preserved despite stricter rules for influential posts. These organic practices demonstrate operational feasibility and indicate latent demand for tiered accountability that transcends any single platform architecture.
We further show that there is a viable regulatory pathway to achieving our proposal. The European Union’s Digital Services Act already requires marketplaces to verify business users and offers a blueprint for identity-linked content duties [61]. Under the UK Online Safety Act, “Category 1” services must give adult users the ability to filter interactions with unverified or anonymous accounts and are required to make identity-verification mechanisms available, without imposing mandatory verification on all users  [62]. By drawing on these precedents, legislators can embed tiered anonymity into safety models and ranking systems without prohibiting pseudonymity outright.
@@
    We chart a jurisdiction-spanning regulatory pathway that leverages existing DSA and Online Safety Act provisions to operationalize the model.
@@ 3 Friction, Identity, and Accountability
The starting point for any meaningful reform of online anonymity must confront a central tension in liberal democracies: the commitment to free expression versus the need to mitigate its weaponization. Classic accounts of speech rights, from John Stuart Mill to modern First Amendment jurisprudence, treat expression as a public good – presumptively beneficial and self-regulating [8, 55, 44]. Yet in algorithmically mediated environments, where LLMs can generate plausible falsehoods at scale and deepfakes blur the boundary between reality and fabrication, virality becomes decoupled from both truth and reputation, and the foundational assumptions underpinning these traditions lose coherence.
Friction – in the form of verification, moderation, or traceability – is often framed as a threat to the open internet [45]. That said, friction is a democratic design feature [88]. In physical communities, social friction arises from reputational consequences, geographic co-presence, and mutual visibility. One is less likely to spread inflammatory falsehoods in a town hall than online, not because one is more moral, but because the social costs are real and immediate. Digital platforms, in contrast, systematically dissolve these frictions. Recommender systems prioritize engagement, not deliberation; speed trumps reflection; and pseudonymity attenuates accountability [39].
@@
We argue that identity obligations must scale with content reach , particularly to mitigate the systemic risks posed by synthetic media as described in Section˜2. This is not a blanket call for real-name policies, which have been rightly criticized for silencing vulnerable speakers [14, 54]. Instead, it is a call for proportional identity calibration [12], wherein pseudonymity is preserved for low-reach users, while higher-tier actors must submit to private identity verification and, ultimately, to structured content review [41]. This approach mirrors how democratic institutions already manage power: with increasing transparency and accountability as influence grows [32].
Our Reddit case study, described in Section˜4, illustrates this principle in practice. As communities expand in size and influence, moderation architectures evolve from permissive to hierarchical: identity checks, posting restrictions, and content approvals become the norm. These organically emergent structures reflect a collective intuition: that scale demands scrutiny and visibility must be earned.
@@ 4 Reddit Case Study in Community Moderation at Scale
Quantitative work finds a positive correlation between subreddit size and the likelihood of (i) entry gates (minimum account age/karma) [31], (ii) pre-publication queues [31, 5], and (iii) ex post identity checks [72]. For example, nearly all large subreddits quietly enforce minimum karma or account-age requirements: smaller communities may require 50 karma, whereas top-tier subreddits often demand well over 1,000 karma points [48]. In some high-visibility subreddits, moderators verify a user’s real identity privately for credibility (e.g. the /r/IAmA forum requires proof of identity for hosts) while allowing the user to remain publicly under a pseudonym. In other words, moderators intuitively impose proportional friction: low-reach users post freely; higher-reach content encounters verification or review. Reddit thus supplies real-world evidence that tiered anonymity is operationally feasible and socially accepted when the costs of influence are borne chiefly by those who wield it.
Table 2: Identity and moderation norms on major social-media platforms as of May 2025. “Tiered” denotes any mechanism in which obligations or scrutiny escalate with audience size or monetization status [^4] Note: Both Facebook and Instagram have limited community moderation in the form of user reports, content-flagging tools, and basic guideline enforcement. However, this does not equate to the Reddit model, where moderation is structurally decentralized: individual subreddits are governed by volunteer moderators with delegated authority to set rules, curate discourse, and shape community norms. Facebook and Instagram lack this layered, self-governing architecture; their “community moderation” is reactive, not constitutive..
@@
Additional scrutiny for 100100k+ channels
@@
Competing platforms provide no comparable venue for community-level rule-making. Facebook real-name enforcement, X’s paid “blue check”, and YouTube’s purely algorithmic filters all exemplify top-down moderation with minimal local discretion or tiering. Comparative studies confirm that Reddit alone relies “more or less on self-moderation by volunteers”, producing a distinctive, multi-layer oversight regime [4]. We summarize our findings regarding identity and moderation norms in all current major social media platforms in Footnote˜4. YouTube provides the closest analogue to our tiered system: “new” channels face automated checks, Partner-Program creators add identity and monetization audits, and six-figure-subscriber channels receive further manual review and provenance badges.
@@ 5 Proposed High-level Technical Implementation
Most large platforms already track granular engagement telemetry (followers, impressions, reshares, watch-time). We propose aggregating these signals into a rolling reach score mapped to tier thresholds proposed in Table˜1. As engagement models differ by platform, each platform must calibrate threshold values to its context. For fairness, reach can be defined in relative terms (e.g. top percentile of active users) to achieve consistency across services. A post that spikes above a threshold would trigger retroactive elevation of the account’s tier. This design ensures that one-off virality is handled without allowing bad actors to repeatedly skirt the limits. Refining the exact reach formula and threshold calibration is left for future work; it will be crucial to balance curbing abuse with not unduly deterring mid-tier creators from legitimate growth.
Platforms should complement hard metrics with contextual triggers such as monetization enrollment or activation of business tools. Precedent exists with Instagram withholding external–link “Swipe-Up” stories until an account reaches $\sim\!10\,000$ followers or holds a business profile, effectively coupling functionality to influence [87]. A similar gating mechanism can enforce tier promotion automatically while minimizing false positives.
@@ 6 Current Legal Precedents and Regulatory Infrastructure
6.1 European Union: Evolving From KYBC Toward Influence-Based Identity Controls
@@
More significantly, Articles 34 and 35 impose systemic risk obligations on Very Large Online Platforms (VLOPs) – defined by a monthly audience threshold – explicitly linking platform reach to responsibility [28]. These risk obligations are particularly salient in the age of generative AI, again linking back to the discussion in Section˜2. This capacity decouples virality from veracity, amplifying the urgency of proportional identity and transparency duties for high-reach users. This sets a critical precedent: the broader a user or platform’s influence, the greater the required diligence. Article 10 further enables identity disclosure in response to illegality, reinforcing a principle of proportionality that mirrors the core logic of tiered anonymity. [22].
Moreover, complementary frameworks like the AI Act and formerly proposed AI Liability Directive further strengthen this trajectory. By requiring labeling of synthetic media and audit trails for AI systems, the EU is already enforcing traceability in high-risk communicative environments [29, 25]. Our proposed tiered framework fits squarely within this expanding digital acquis. These instruments, taken together, suggest that scalable identity obligations based on content reach are not only compatible with EU law – they are its logical extension.
@@ 6.2 United Kingdom: The Online Safety Act and Voluntary Verification
In contrast, the UK’s Online Safety Act 2023 [62] establishes a statutory duty of care on digital platforms, particularly those classified as Category 1 services – platforms with significant reach and functionality. Under the accompanying Categorization of Regulated Services (Threshold Conditions) Regulations 2024 [79], these platforms are required to offer adult users the option to verify their identity and to provide tools enabling content filtering based on verification status. This framework embeds optional identity verification with a reputational signalling layer, while preserving the right to anonymity, laying the conceptual groundwork for our proposed tiered identity regime.
However, this identity framework remains voluntary and reputational rather than mandatory and enforceable. Users may choose to verify themselves, and others may opt to filter content accordingly; but no binding obligations are imposed on high-reach users who remain anonymous. Existing legal tools expose both the possibilities and the limits of the current system. Norwich Pharmacal Orders can compel disclosure only when platforms already hold identity data, a weak point given that many accounts have never been verified. [78] Similarly, the UK–US CLOUD Act facilitates law-enforcement access to stored data but is not an identity-disclosure regime and does not ensure that platforms possess verified identities.[77] These mechanisms show that anonymity is constrained in principle yet remain inadequate for ensuring accountability among high-reach anonymous users.
The current UK regulatory landscape lacks a proactive mechanism linking user influence – measured by visibility, engagement, or monetization – to identity obligations. We argue that this omission is increasingly untenable in an era of algorithmic virality where individuals can rapidly attain significant reach with little to no accountability.
A logical evolution of the Online Safety Act would be to mandate identity verification for users who exceed a defined influence threshold, as discussed in Section˜5. Such a reform would convert identity verification from a reputational indicator into a mechanism of enforceable accountability. By embedding this obligation within the existing statutory framework, the UK could pioneer a rights-preserving yet responsibility-tiered model of online governance – one that maintains anonymity for everyday users while ensuring that high-reach actors meet proportionate standards of transparency and legal traceability.
@@ 6.3 United States: First-Amendment Boundaries and Conditional Immunity
The United States presents the most challenging jurisdiction for any form of compelled identity regulation due to robust First Amendment protections and the shield of Section 230555Section 230 provides that online platforms are not treated as the “publisher” of user content, effectively immunizing them from most civil liability for third-party posts of the Communications Decency Act [75]. American courts have repeatedly upheld the right to anonymous speech, particularly in digital spaces. Landmark cases such as Doe v. Cahill and Dendrite Int’l, Inc. v. Doe No. 3 [19] require plaintiffs seeking to unmask anonymous users to meet stringent standards, such as presenting a prima facie case of harm and passing a balancing test that weighs the speaker’s right to anonymity.
Despite this, momentum is growing at the federal level toward rethinking the blanket nature of Section 230 immunity. Legislative proposals, including bipartisan efforts, have increasingly considered conditioning immunity on a platform’s compliance with transparency and good-faith content moderation practices [70, 21]. Crucially, however, the constitutional limits on compelled identity disclosure derive from the First Amendment, not from Section 230. Section 230 protects platforms from liability for third-party content; it does not shield them from disclosure obligations or other regulatory requirements, which are instead constrained by First Amendment doctrine. Recent legislative steps also target AI-driven harms directly – for example, the 2025 TAKE IT DOWN Act criminalizes the distribution of non-consensual intimate imagery (including AI-generated deepfakes) and mandates swift removal of such content [76]. Rather than mandating universal real-name usage (which would invite constitutional challenges), these proposals suggest a path for indirect, incentive-based regulation that respects constitutional limits while introducing mechanisms of accountability.
In this context, our proposed tiered identity approach offers a legally viable and technically feasible approach. Platforms could retain full Section 230 protections only if they implement a reach-based anonymity tier system. Basing the framework on influence, or other platform-side metrics such as engagement telemetry and monetization enrolment, allows the system to remain content-neutral, which is crucial for surviving constitutional scrutiny [38, 20]
@@ 8 The Global Momentum for Conditional Pseudonymity
The international policy environment is already increasingly converging around the idea that identity obligations should scale with user influence. Among the first state efforts to discipline online anonymity was South Korea’s real-name verification regime, a sweeping identity-disclosure mandate intended to eliminate anonymous participation across major platforms. That approach proved both legally unsustainable and practically ineffective. The Korean Constitutional Court invalidated the policy for violating freedom of expression, and subsequent research showed it failed to reduce online harms in any measurable way [60, 52]. The lesson was clear: blanket identity mandates are blunt instruments that overreach without precision.
Since then, regulatory energy has shifted toward more granular, influence-sensitive models. In India, the 2023 Draft Digital India Bill introduces a risk-based classification framework for digital intermediaries, suggesting a shift toward more nuanced regulatory obligations based on the type and scale of service; without explicitly extending these obligations to individual users or calibrating them to user influence [1]. Australia’s eSafety Commissioner has advanced similar proposals, calling for the traceability of high impact accounts, particularly those linked to harmful or AI-generated content [24]. Meanwhile, the European Commission has initiated consultations on “influence transparency”, exploring how verification requirements might apply to accounts disseminating politically sensitive or synthetic media [61, 26].
@@
9 Enforcement Challenges and Safeguards
Implementing a tiered anonymity framework in a globally interconnected internet ecosystem presents significant enforcement challenges – large commercial platforms operating across jurisdictions may resist added verification burdens, and users might evade restrictions by migrating to less-regulated services or using technical workarounds (e.g. bots to artificially reset reach). While Reddit shows that moderation hierarchies can emerge organically, its reliance on volunteer governance is difficult to replicate on more commercial platforms like Meta, YouTube, or X. These platforms operate across multiple jurisdictions but often default to the legal norms of their home country – typically the United States – resulting in fragmented regulatory oversight [7].
To scale tiered anonymity, enforcement must be institutional, driven by governments and platforms rather than individual users. This coordinated institutional response is essential: voluntary self-regulation by platforms alone has thus far failed to curb disinformation, prioritizing profits over safety – “fundamentally self-regulation by social media platforms has failed to achieve the promised public policy benefit of improving the quality of the information ecosystem” [42]. Governments in regions such as the EU, UK, and US already exercise regulatory authority over platforms operating within their borders. This authority can be extended extraterritorially, as seen with the EU’s General Data Protection Regulation (GDPR), which extends obligations beyond EU borders through data adequacy requirements and reputational enforcement mechanisms [65, 3].